# Non-Iterative Coordination of Interconnected Power Grids via Dimension-Decomposition-Based Flexibility Aggregation

### v2

@@ Abstract
The bulk power grid is divided into regional grids interconnected with multiple tie-lines for efficient operation. Since interconnected power grids are operated by different control centers, it is a challenging task to realize coordinated dispatch of multiple regional grids. A viable solution is to compute a flexibility aggregation model for each regional power grid, then optimize the tie-line schedule using the aggregated models to implement non-iterative coordinated dispatch. However, challenges such as intricate interdependencies and curse of dimensionality persist in computing the aggregated models in high-dimensional space. Existing methods like Fourier-Motzkin elimination, vertex search, and multi-parameter programming are limited by dimensionality and conservatism, hindering practical application. This paper presents a novel dimension-decomposition-based flexibility aggregation algorithm for calculating the aggregated models of multiple regional power grids, enabling non-iterative coordination in large-scale interconnected systems. Compared to existing methods, the proposed approach yields a significantly less conservative flexibility region. The derived flexibility aggregation model for each regional power grid has a well-defined physical counterpart, which facilitates intuitive analysis of multi-port regional power grids and provides valuable insights into their internal resource endowments. Numerical tests validate the feasibility of the aggregated model and demonstrate its accuracy in coordinating interconnected power grids.
@@
$\overline{(\cdot)},\underline{(\cdot)}$ The upper and lower limits parameters of variables.
@@
$\operatorname{space}(\bm{x})$ The space where the vector $\bm{x}$ resides.
@@ I-A Background and Motivation
The modern power grids in China, North America and Europe are all interconnected systems. To facilitate administration and coordination, the bulk power grid is segmented into several regional power grids (RPGs) based on geographical areas, interconnected by AC and DC tie-lines [1]. The proportion of renewable energy and the varying weather conditions in various regions create complementary patterns in peak electricity supply and demand. Energy transmission through AC and DC tie-lines provides cross-regional energy support ensures stable and economic grid operations of the whole system.
@@
However, achieving optimal coordinated dispatch across multiple RPGs presents significant challenges, as each RPG operates under its own control center. Data privacy concerns and variations in data models further complicate the coordination between these grids. In addition, the sheer volume of devices and network data makes centralized management and optimization of the entire bulk power grid impractical. A commonly adopted approach is to coordinate RPGs using a decentralized optimization framework, then the optimal dispatch problem of each RPG is treated as a sub-problem, with the overall coordination of the bulk power grid achieved through an iterative process. A novel decomposition method with quasi-second order convergence rate proposed by [2] generalizes the application scope from linear to generalized nonlinear problems, which is a substantial advancement in decomposition techniques that significantly reduces the iteration number and computation burden among control centers. Nevertheless, guaranteeing convergence in such iterative schemes remains challenging, and diagnosing and remedying divergence when it arises can be difficult. Additionally, coordinated dispatch requires synchronized, real-time communication among all regional control centers, which creates significant obstacles for practical engineering implementation.
@@
An effective solution to these challenges is to calculate the power transmission flexibility region of tie-lines between RPGs, which is also called tie-line security region in [3, 4]. Then the coordination of interconnected power grids can be achieved without the need for iterative coordination [5]. A schematic diagram illustrating the aggregation of regional power grids is shown in Fig. 1.
@@
Figure 1: Aggregate the regional power grids for coordination of interconnected systems.
@@
By employing an aggregated model, each RPG can focus solely on the external characteristics of tie-lines, without considering complex internal network constraints, thereby significantly reducing the complexity of the optimal coordination of the bulk grid. The flexibility aggregation results can also be used as a comprehensive assessment of the overall RPG, providing valuable insights for the resource endowment inside. At the same time, device data and information within each RPG is effectively protected through aggregation.
@@
Calculating the aggregated model of an RPG still faces major challenges. Firstly, the decision variables in an RPG are numerous. All these variables are coupled with each other by large-scale spatio-temporal coupling constraints, including the network constraints at single time slices, and the temporal coupling constraints of generation units across all time slots. All the variables and the coupling constraints form a very complex high-dimensional space. Besides, an RPG connects to other RPGs through multiple AC and DC tie-lines, which are coupled with each other due to the network constraints. It is also difficult to explicitly express these spatio-temporal coupling relationships as constraints among these tie-lines.
@@
Some existing studies have proposed several types of methodologies for computing the flexibility aggregation model of a system. The Fourier-Motzkin elimination method is proposed in [6, 7] to calculate the flexibility set of power system by eliminating the internal variables of the system. The concept of dispatchable region is presented in [8] and calculated by iteratively creating boundary hyperplanes of the feasible region. The progressive vertex enumeration method is proposed in [9] to calculate the projected flexibility region of power system by iteratively searching the extreme points. Besides, the work in [3] presents an evaluation method for the secure region of a single tie-line between regional power grids. The multi-parameter programming based algorithms are presented in [10] and its enhanced acceleration algorithm is proposed in [11]. Security region of renewable energy integration is calculated in [12] with multi-parameter programming. Geometry-based methods also offer an approach to calculate the aggregation model, where the aggregated flexibility region is approximated using various geometric shapes, such as cubes [13], ellipsoids [14], and polytopes shaped by storage and/or generator constraints [15, 16, 17]. Moreover, to deal with the high-dimension coupling among multiple tie-lines in an RPG, [4] decomposes the flexibility region along the temporal dimension and recombines the lower-dimensional regions with Cartesian product.
@@
In practical power grids, the presence of multiple tie-lines between regional grids creates a multi-port system, significantly increasing computational complexity. Existing methods such as Fourier-Motzkin elimination, vertex search, multi-parameter programming, and direct geometric projection all suffer from the curse of dimensionality. As a result, these approaches often fail to compute the flexibility region of RPGs within a limited time, limiting their practical application in engineering. Furthermore, existing methods exhibit excessive conservatism, producing flexibility regions that are much smaller than the actual regions. This limitation prevents them from achieving optimal solutions in dispatch applications and hinders their ability to attain economic optimality.
I-B Contributions and Paper Organization
@@ II-A Physical and Mathematical Interpretation for the Flexibility Aggregation Problem
From a physical perspective, a system consists of numerous interconnected internal devices. These devices influence each other and interact with the external environment through boundary devices, such as feeders or tie-lines, enabling energy and information exchange. To aggregate the flexibility of the entire energy system, the internal devices and coupling networks are compressed into a flexible equivalent aggregation model. Then, the system’s overall flexibility, viewed externally, can be represented solely by the boundary devices. The internal structure of this system is treated as a black box [20].
Considering the mathematical model of all devices in the system, the system is governed by the operational and security constraints involving the internal variables $\bm{x}^{\text{int}}\in\mathbb{R}^{m}$ and the boundary variables $\bm{x}^{\text{bd}}\in\mathbb{R}^{n}$. If these constraints can be approximated or relaxed in a linear form, they constitute the feasible region $\mathcal{R}_{\text{H}}$ of the whole system as follows:
@@
Since the dimensionality reduction projection of a polytope can retain the form of polytope, the projected polytope can be expressed as
@@ II-B Overview of Geometric-Based Methods
Using the physical properties and geometric interpretation inherent to flexibility aggregation, shape template-based methods can be used to approximate the flexibility region $\mathcal{R}_{\text{L}}$ [16]. The main physical properties that can be used are the flexibility characteristics of the internal devices. Since the system’s aggregated flexibility is determined by the combination of all devices it contains, if the flexibility regions of these devices share similar shapes, this similarity can be used as a shape template to represent the system’s overall flexibility.
With the linearized approximation [23, 24] of the flexibility regions of devices, they can be expressed in the common linear form:
$\bm{A}\bm{x}_{i}\leq\bm{b}_{i}$ (4)
where $\bm{A}$ is a parameter matrix shared by all devices with the same flexibility property, representing the common shape of their flexibility regions. $\bm{x}_{i}$ denotes the decision variables of the $i$-th device, and $\bm{b}_{i}$ refers to the constraint parameters of the $i$-th device. Then, this expression similarity can be used as a shape template for the inner-approximated region as follows:
^$\hat{\mathcal{R}}_{\text{L}}(\bm{b})=\left\{\bm{x}^{\text{bd}}\left\lvert\,\bm{A}\bm{x}^{\text{bd}}\leq\bm{b}\right.\right\},\hat{\mathcal{R}}_{\text{L}}\subset\mathbb{R}^{n}$ (5)
where $\bm{A}$ is the same matrix as in (4). The vector $\bm{b}$ represents the parameters of the inner-approximated region ^$\hat{\mathcal{R}}_{\text{L}}$. These parameters need to be adjusted to ensure that the condition ^$\hat{\mathcal{R}}_{\text{L}}(\bm{b})\subseteq\mathcal{R}_{\text{L}}$ is satisfied. This adjustment can be achieved using our previously developed polytope-bound shrinking algorithm, as detailed in [16, 20].
@@
Fig. 4 is an illustrative diagram that explains the procedure of the algorithm, with the detailed algorithms available in [16]. Inspired by geometric perspective, this algorithm starts from a circumscribed polytope of $\mathcal{R}_{\text{L}}$ with the given template shape and gradually shrinks the boundaries of the inner-approximated region ^$\hat{\mathcal{R}}_{\text{L}}(\bm{b}_{(k)})$ by gradually adjusting the values of parameters $\bm{b}_{(k)}$ in the $k$-th iteration. Additionally, if all nonzero elements in each row of the matrix $\bm{A}$ in (5) have the same sign, the modified bound-shrinking method [22] can be applied to accelerate calculations.
For the cases where there is no explicit mathematical expression like (5) of the flexibility region—due to complex internal device variables or intricate coupling relationships among boundary variables—the flexibility region can instead be inner-approximated using a high-dimensional ellipsoid shape template [21] as (6).
@@
To achieve the best possible approximation, the volume of the high-dimensional ellipsoid should be maximized. This process is analogous to inflating an ellipsoid like a balloon within the projected region until it touches the boundaries, reaching its maximum volume.
This problem can be expressed as the following robust optimization problem:
@@
Due to the difficulty in solving the optimization problem (7) in high-dimensional space, the problem can be simplified by introducing specific decision policies between variables $\bm{\xi}$ and $\bm{x}^{\text{int}}$. In engineering applications, adding linear or quadratic policies to the robust problem (7) is a common practice, which allows for an equivalent reformulation as a semi-definite programming (SDP) problem. Details on calculating the parameters $\bm{E}$ and $\bm{e}$ using an equivalent SDP can be found in [21].
III Mathematic Model of a Regional Power Grid
@@ III-A Network and Power Balance Constraints
The operational constraints in a regional power grid are as follows
@@
where (8a) calculates the net injection power $\bm{p}^{\text{inj}}_{t}$ of all buses at time $t$. The matrices $\bm{A}^{\text{tpg}}$, $\bm{A}^{\text{rgu}}$, and $\bm{A}^{\text{tie}}$ represent device-bus associations, mapping the indices of each respective device to the corresponding bus indices. (8b) is the transmission line power capability constraints. $\bm{S}^{\text{PTDF}}$ denotes the power transfer distribution factor (PTDF) matrix. The voltage angles of the boundary buses that are directly connected to AC tie-lines, denoted as $\bm{\theta}_{t}^{\text{bd}}$, are calculated in (8c). $\theta_{t}^{\text{ref}}$ denotes the voltage phase angle of the reference node in this RPG relative to the global reference node. $\bm{B}$ denotes the susceptance matrix. $\bm{M}^{\text{bd}}$ is used to select all the indices of boundary buses that directly connected to AC tie-lines

### v3

@@ Abstract
The bulk power grid is divided into regional grids interconnected with multiple tie-lines for efficient operation. Since interconnected power grids are operated by different control centers, coordinating dispatch across multiple regional grids is challenging. A viable solution is to compute a flexibility aggregation model for each regional power grid; then optimize the tie-line schedule using the aggregated models to implement non-iterative coordination. However, challenges such as intricate interdependencies and the curse of dimensionality persist in computing the aggregated models in high-dimensional space. Existing methods like Fourier-Motzkin elimination, vertex search, and multi-parameter programming are limited by dimensionality and conservatism, hindering their practical application. This paper presents a novel dimension-decomposition-based flexibility aggregation algorithm for calculating the aggregated models of multiple regional power grids, enabling non-iterative coordination in large-scale interconnected systems. Compared to existing methods, the proposed approach yields a significantly lower computational complexity and a less conservative flexibility region. The derived flexibility aggregation model for each regional power grid has a well-defined physical counterpart, which facilitates intuitive analysis of multi-port regional power grids and provides valuable insights into their internal resource endowments. Numerical tests validate the feasibility of the aggregated model and demonstrate its precision in coordinating interconnected power grids.
@@
$\bm{p}^{\text{ld}}_{t}$ Vector composed of buses’ load power at time $t$.
@@
$\overline{(\cdot)},\underline{(\cdot)}$ The upper- and lower-bound parameters of variables.
@@
$\operatorname{space}(\bm{x})$ The space where the vector $\bm{x}$ resides. $\operatorname{space}(\bm{x})=\mathbb{R}^{n}$, where $n:=\operatorname{len}(\bm{x})$
@@ I-A Background and Motivation
The modern power grids in China, North America and Europe are all interconnected systems. To facilitate administration and coordination, the bulk power grid is segmented into several regional power grids (RPGs) based on geographical areas and interconnected by AC and DC tie-lines [1]. The proportion of renewable energy and the varying weather conditions in various regions create complementary patterns in peak electricity supply and demand. Energy transmission through AC and DC tie-lines provides cross-regional energy support and ensures stable and economical operation of the whole system.
@@
However, achieving optimal coordinated dispatch across multiple RPGs presents significant challenges, as each RPG operates under its own control center. Data privacy concerns and variations in data models further complicate the coordination between these grids. In addition, the sheer volume of devices and network data makes centralized management and optimization of the entire bulk power grid impractical. A commonly adopted approach is to coordinate RPGs using a decentralized optimization framework; the optimal dispatch problem of each RPG is then treated as a sub-problem, with the overall coordination of the bulk power grid achieved through an iterative process. An iteration-based decomposition method, with quasi-second-order convergence rate, proposed in [2], generalizes from linear to non-linear problems. This represents a substantial advancement in decomposition techniques, significantly reducing the number of iterations and computational burden at each control center. Nevertheless, guaranteeing convergence in such iterative schemes remains challenging, and diagnosing and remedying divergence when it arises can be difficult. Additionally, coordinated dispatch requires synchronized real-time communication between all regional control centers, which creates significant obstacles in practical engineering implementation.
@@
An effective solution to these challenges is to calculate the power transmission flexibility region of tie-lines between RPGs, which is also called tie-line security region in [3, 4]. Then coordination can be achieved via a non-iterative approach [5]. A schematic diagram illustrating the aggregation of regional power grids is shown in Fig. 1.
@@
Figure 1: Aggregation of regional power grids for coordinated dispatch of interconnected systems.
@@
By employing an aggregated model, each RPG can focus solely on the external characteristics of tie-lines, without considering complex internal network constraints, thereby significantly reducing the complexity of the optimal coordination of the bulk grid. The results of flexibility aggregation can also be used as a comprehensive assessment of the overall RPG, providing valuable information for the endowment of resources within the system [6]. At the same time, device data and information within each RPG is effectively protected through aggregation.
A comprehensive comparison between distributed iterative and aggregated model-based non-iterative methods is summarized in TABLE I.
TABLE I: Comparison Between Distributed Iterative and Aggregated Model-Based Non-Iterative Methods
Characteristic Iterative Non-Iterative
Convergence Not guaranteed; failure causes are often unclear Guaranteed in a single step
Iteration count Multiple (tens to hundreds) One-time computation
Communication Requires real-time, synchronous updates per iteration; vulnerable to bottlenecks from the slowest participants One-time aggregated model exchange
Real-time coordination computational burden High cumulative runtime Moderate one-time overhead
Scalability Iterations increase with the number of sub-problems Highly scalable due to decoupling
Implementation Requires complex coordination protocols Simple and modular deployment
I-B Literature Review
@@
Calculating an RPG’s aggregated model still presents major challenges. Firstly, the decision variables in an RPG are numerous. All these variables are coupled with each other by large-scale spatio-temporal coupling constraints, including the network constraints at single time slices and the temporal coupling constraints of generation units across all time slots. All the variables and the coupling constraints form a very complex, high-dimensional space. In addition, an RPG connects to other RPGs through multiple AC and DC tie-lines, which network constraints further interconnect. It is also difficult to explicitly express these spatio-temporal coupling relationships as constraints among these tie-lines.
@@
Some existing studies have proposed several types of methodologies for computing the flexibility aggregation model of a system. The Fourier-Motzkin elimination method is proposed in [7, 8] to calculate the flexibility region of the power system by eliminating the internal variables of the system. The concept of the dispatchable region is presented in [9] and calculated by iteratively creating boundary hyperplanes of the feasible region. The progressive vertex enumeration method is proposed in [10] to calculate the projected flexibility region of the power system by iteratively searching for the extreme points. In addition, the work in [3] presents an evaluation method for the secure region of a single tie-line between regional power grids. The multi-parameter programming-based algorithms are presented in [11], and an enhanced acceleration scheme is proposed in [12]. The security region of renewable energy integration is calculated in [13] with multi-parameter programming. Geometry-based methods also offer an approach to calculate the aggregation model, where the aggregated flexibility region is approximated using various geometric shapes, such as cubes [14], ellipsoids [15], and polytopes shaped by storage and/or generator constraints [16, 17, 18]. Moreover, to deal with the high-dimensional coupling among multiple tie-lines in an RPG, [4] decomposes the flexibility region along the temporal dimension and recombines the lower-dimensional regions with the Cartesian product.
@@
In practical power grids, the presence of multiple tie-lines between regional grids creates a multi-port system, significantly increasing computational complexity. Existing methods such as Fourier-Motzkin elimination, vertex search, multi-parameter programming, and direct geometric projection all suffer from the curse of dimensionality. As a result, these approaches often fail to compute the flexibility region of RPGs in a limited time, limiting their practical application in engineering. Furthermore, existing methods exhibit excessive conservatism, producing flexibility regions that are much smaller than the actual regions. This limitation prevents them from achieving optimal solutions in dispatch applications and hinders their ability to attain economic optimality.
I-C Contributions and Paper Organization
@@ II-A Physical and Mathematical Interpretation for the Flexibility Aggregation Problem
From a physical perspective, a system consists of numerous interconnected internal devices. These devices influence each other and interact with the external environment through boundary devices, such as feeders or tie-lines, enabling energy and information exchange. To aggregate the flexibility of the entire energy system, the internal devices and coupling networks are compressed into a flexible equivalent aggregation model. Then, the system’s overall flexibility, viewed externally, can be represented solely by the boundary devices. The internal structure of this system is treated as a black box.
Considering the mathematical model of all the devices in the system, the system is governed by the operational and security constraints involving the internal variables $\bm{x}^{\text{int}}\in\mathbb{R}^{m}$ and the boundary variables $\bm{x}^{\text{bd}}\in\mathbb{R}^{n}$. If these constraints can be approximated or relaxed in a linear form, they constitute the feasible region $\mathcal{R}_{\text{H}}$ of the whole system as follows:
@@
Since the dimensionality reduction projection of a polytope can retain the form of a polytope, the projected polytope can be expressed as
@@ II-B Overview of Geometric-Based Methods
Using the physical properties and geometric interpretations inherent in flexibility aggregation, shape template-based methods can effectively approximate the flexibility region $\mathcal{R}_{\text{L}}$ [17]. The key physical properties utilized in this context are the flexibility characteristics of the constituent internal devices. Recent geometry-based methods typically employ polytopal and ellipsoidal templates, derived from physical property analysis, to inner-approximate the aggregated flexibility region $\mathcal{R}_{\text{L}}$.
Since the system-level aggregated flexibility emerges from the collective behavior of all internal devices, if their individual flexibility regions share certain structural properties, the system’s flexibility region can be approximated using a polytopal shape template defined by explicit linear constraints. For instance, if all internal devices of an energy system exhibit energy storage property, such as energy storage devices, electric vehicles, and thermal controllable loads, the system’s flexibility can be aggregated into an equivalent energy storage model. That is,
$\underline{p}_{t}\leq p_{t}\leq\overline{p}_{t},\forall 1\leq t\leq T$ (4a)
$\frac{\underline{E}_{t_{1},t_{2}}}{\Delta t}\leq\sum_{\tau=t_{1}}^{t_{2}}p_{\tau}\leq\frac{\overline{E}_{t_{1},t_{2}}}{\Delta t},\forall 1\leq t_{1}<t_{2}\leq T$ (4b)
The shape template for the inner-approximated region can be expressed in a compact matrix form as follows:
^$\hat{\mathcal{R}}_{\text{L}}(\bm{b})=\left\{\bm{x}^{\text{bd}}\left\lvert\,\bm{A}\bm{x}^{\text{bd}}\leq\bm{b}\right.\right\}$ (5)
where the boundary variable vector $\bm{x}^{\text{bd}}$ is composed of $p_{t}$ across all time slots in this case. The matrix $\bm{A}$ is the parameters that correspond to the constraints (4), expressed in the storage form. The vector $\bm{b}$ represents the parameters of the inner-approximated region ^$\hat{\mathcal{R}}_{\text{L}}$, composed of all the paremeters $\underline{p}_{t}$, $\overline{p}_{t}$, $\nicefrac{{\underline{E}_{t_{1},t_{2}}}}{{\Delta t}}$, and $\nicefrac{{\overline{E}_{t_{1},t_{2}}}}{{\Delta t}}$ in this case. The parameters in $\bm{b}$ need to be adjusted to ensure that condition ^$\hat{\mathcal{R}}_{\text{L}}(\bm{b})\subseteq\mathcal{R}_{\text{L}}$ is satisfied. This adjustment can be achieved using our previously developed polytope-bound shrinking algorithm, as detailed in [17].
@@
Fig. 4 is an illustrative diagram that explains the procedure of the algorithm, with the detailed algorithms available in [17]. Inspired by geometric perspective, this algorithm starts from a circumscribed polytope of $\mathcal{R}_{\text{L}}$ with the given template shape and gradually shrinks the boundaries of the inner-approximated region ^$\hat{\mathcal{R}}_{\text{L}}(\bm{b}_{(k)})$ by gradually adjusting the values of parameters $\bm{b}_{(k)}$ in the $k$-th iteration. Furthermore, if all nonzero elements in each row of the matrix $\bm{A}$ in (5) have the same sign, the modified bound shrinking method [22] can be applied to accelerate calculations.
When the flexibility region lacks an explicit constraint expression like (5), due to the presence of complex internal device variables or intricate coupling among boundary variables, it can be inner-approximated using a high-dimensional ellipsoid shape template [21], as illustrated in (6).
@@
To achieve the best possible approximation, the volume of the high-dimensional ellipsoid should be maximized. This process is analogous to inflating an ellipsoid like a balloon within the projected region until it touches the boundaries and reaches its maximum volume. It can be expressed as the following robust optimization problem:
@@
where $\bm{x}^{\text{bd}}:=\bm{E}\bm{\xi}+\bm{e}$. Due to the difficulty in solving (7) in high-dimensional space, the problem can be simplified by introducing specific decision policies between variables $\bm{\xi}$ and $\bm{x}^{\text{int}}$. In engineering applications, adding linear or quadratic policies to the robust problem (7) is a common practice, which allows for an equivalent reformulation as a semi-definite programming (SDP) problem. Details on calculating the parameters $\bm{E}$ and $\bm{e}$ using an equivalent SDP can be found in [21].
In summary, geometric-based methods commonly employ polytopal and ellipsoidal shape templates to inner-approximate the aggregated flexibility region $\mathcal{R}_{\text{L}}$. A template for polytopal shapes that match the flexibility properties can more accurately capture the boundaries of the actual flexibility region, thereby reducing conservativeness. On the other hand, ellipsoidal templates are well-suited for representing irregular flexibility regions with few parameters. However, their smooth and symmetric nature may result in more conservative approximations. In practical applications, the choice between these templates can be guided by the availability of explicit constraints and the characteristics of the constraints within the system.
III Mathematical Model of a Regional Power Grid
@@ III-A Network and Power Balance Constraints
The operational constraints in a regional power grid are as follows.
@@
where (8a) calculates the net injection power $\bm{p}^{\text{inj}}_{t}$ of all buses at time $t$. $\bm{p}^{\text{tpg}}_{t}$, $\bm{p}^{\text{rgu}}_{t}$ are vectors composed of thermal power generators and renewable generators’ output power at time $t$. $\bm{p}^{\text{tie}}_{t}$ denotes the vector composed of tie-lines’ output power at time $t$. $\bm{p}^{\text{ld}}_{t}$ denotes the vector composed of buses’ load power at time $t$. The matrices $\bm{A}^{\text{tpg}}$, $\bm{A}^{\text{rgu}}$, and $\bm{A}^{\text{tie}}$ represent device-bus associations, mapping the indices of each respective device to the corresponding bus indices. (8b) is the transmission line power capability constraints. $\bm{S}^{\text{PTDF}}$ denotes the power transfer distribution factor (PTDF) matrix. The voltage angles of the boundary buses that are directly connected to AC tie-lines, denoted as $\bm{\theta}_{t}^{\text{bd}}$, are calculated in (8c). $\theta_{t}^{\text{ref}}$ denotes the voltage phase angle of the reference node in this RPG relative to the global reference node. $\bm{B}$ denotes the susceptance matrix. $\bm{M}^{\text{bd}}$ is used to select all the boundary buses indices that are directly connected to AC tie-lines.
Remark: The regional power grid is modeled using the DC power flow formulation in (8), with a focus solely on active power variables. To further incorporate reactive power, the model can be generalized to a high-precision linearized AC power flow framework using the methods proposed in [23, 24]. The flexibility aggregation method proposed herein remains valid as long as linear relationships between variables are maintained.